OpenStudy (anonymous):

(6.3 x 10^-2)/(2.1 x 10^4) solve in scientific notation

OpenStudy (anonymous):

.0063 / 21000 Now just divide it

OpenStudy (anonymous):

Oops, I mean .063, not .0063

OpenStudy (anonymous):

So it is .063 / 21000

OpenStudy (unklerhaukus):

\[\frac{6.3 \times 10^{-2}}{2.1 \times 10^4} =\frac{6.3}{2.1}\times10^{-2-4}\]

OpenStudy (anonymous):

Thank you!